OpenStudy (anonymous):

solve x^2/3 - (25/49)=0

OpenStudy (anonymous):

Are you looking for exact or approximate answers?

OpenStudy (anonymous):

im not sure. it just says solve the following equation. i would guess exact

OpenStudy (whpalmer4):

\[\frac{x^2}{3} - \frac{25}{49} = 0\] \[\frac{x^2}{3} = \frac{25}{49} = \frac{5^2}{7^2} = (\frac{5}{7})^2\] Can you solve that for x? Remember, there will be a negative and a positive root.

OpenStudy (anonymous):

im sorry i wrote it wrong the x^2/3 should be x^(2/3)

OpenStudy (whpalmer4):

\[x^{2/3} - \frac{25}{49} = 0\] is the real problem?

OpenStudy (anonymous):

correct

OpenStudy (whpalmer4):

That isn't too hard. Move the fraction to the right hand side, and then undo that gnarly exponent. Remember, \(x^{\frac{2}{3}} = \sqrt[3]{x^2}\).

OpenStudy (anonymous):

so now i have x^2= (25/49)^3

OpenStudy (whpalmer4):

Yes. Take the square root of the right hand fraction before cubing it, that will be easier!

OpenStudy (whpalmer4):

Do you have a final answer?

OpenStudy (anonymous):

no haha i have no idea why take the sqrrt of the right side